OpenStudy (mackenzie2013):

How many times will 1/14 go into 6 7/10? Please explain how to get answer :)

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

\[6\tfrac{7}{10}\div \frac{1}{14}\] better to turn in to an improper fraction \[\frac{67}{10}\div\frac{1}{14}=\frac{67}{10}\times \frac{14}{1}\]